A bond currently has a price of $1,050 and is selling to yield 5%. If the yield increases 23bps, the price of the bond will go down to $1,023. This implies that the:

a) Bond's Modified Duration =

b) Bond's Duration =
Answer #1

Modified Duration=-% change in bond price/change in yield=-(1023/1050-1)/(0.23%)=11.18012

Duration=11.18012*(1+5%)=11.73913